OpenStudy (anonymous):

find the inverse of the following function. find the domain range and asymptotes of each function f(x)= in(8-x)

OpenStudy (jdoe0001):

f(x)= in(8-x)? f(x)= ln(8-x)

OpenStudy (anonymous):